OpenStudy (anonymous):

[12T^3/6s^4 ]cubed simplify expression

OpenStudy (aakashsudhakar):

When you cube a fraction with variables in the numerator and denominator, it's as simple as cubing each expression separately. So let's do this individually. (12T^3)^3 = 1728T^9 (6s^4)^3 = 216s^12 So the fraction becomes [(1728T^9)/(216s^12)] Fortunately, 216 is actually a factor of 1728. So divide and simplify. 1728/216 = 8 The fraction becomes [(8T^9)/(s^12)].

OpenStudy (anonymous):

thanks i thought that was the way to solve it